OpenStudy (anonymous):

use Gauss's approach to find the following sums (do not use formulas). 1+3+5+7+...+997 what is the sum of the sequence?

OpenStudy (anonymous):

thisis realy confusing me its like i almost keep understanding it then i dont understand and get confused!

OpenStudy (anonymous):

gauss a approach is to add up forwards and backwards \[S=1+3+5+...+997\] \[S=997 + 995 + ... + 3 + 1\] \[2S=998 998 + ... + 998= 499\times 998\] \[ S=499^2\]

OpenStudy (anonymous):

ok that was a typo, second line should have been \[S= 998 + 998 + ... + 998=499\times 998\] but this is a silly way to do this

OpenStudy (anonymous):

1+3+5++.....+997 997+995+993+....+1 =998+998+......+998(499 times) so =998*499/2

OpenStudy (anonymous):

it is always true that \[\sum_{k=1}^n(2k-1)=n^2\]

OpenStudy (anonymous):

where did the 499 come from??

OpenStudy (anonymous):

good question is how do we know 499

OpenStudy (anonymous):

oh right, you just asked it !

OpenStudy (anonymous):

all the numbers are odd, so they all look like \[2k-1\] for k = 1,2,3,... and you want to know how many terms there are, because that is what you have to multiply by. set \[2k-1=997\] \[2k=998\] \[k=998\div 2=499\]

OpenStudy (anonymous):

omg thats so confusing!!

OpenStudy (anonymous):

however, it should be entirely clear that the sum the first n odd numbers is n^2 \[1=1^2\] \[1+3=2^2\] \[1+3+5=3^2\] \[1+3+5+7=4^2\] \[1+3+5+7=5^2\] etc

OpenStudy (anonymous):

can i give u another one and show me again with more detail steps

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

48+49+50+51+... +133

OpenStudy (anonymous):

but basically it comes down to figuring out how many terms you have. and although summations can be confusing i am sure you are not confused by how to solve \[2k-1=997\] for k

OpenStudy (anonymous):

again using gauss?

OpenStudy (anonymous):

is their an easier way to solve these without so many steps?

OpenStudy (anonymous):

yes same process

OpenStudy (anonymous):

it is as the last example. one half the number of terms times (first plus last term)

OpenStudy (anonymous):

i really need to understnad this by today cuz this assignment was due yesterday and its so long and i still didnt finish it i have to submit it today!

OpenStudy (anonymous):

in the last example the first term is 48, last term is 133 and their sum is \[48+133=181\] now how many terms are there? should get 86 terms. so your answer will be \[\frac{1}{2}\times 86\times 181\]

OpenStudy (anonymous):

so 86x181/2 ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

its like it makes sense but at the same time i dont get it lol

OpenStudy (anonymous):

so what do you need all together to solve these additions : first term \[a_1\] last term \[a_n\] number of terms. \[n\] then it will be \[S =\frac{n}{2}(a_1+a_n)\]

OpenStudy (anonymous):

you will get it if you do a bunch of examples